a wire of resistance 5 ohm is drawn out to its new length, three times its original length. the resistance of new wire is..

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

R =þ * l / A

Volume is constant

l *A = 3 l * A/3

So

5 =þ l/A
Rnew = 9 * 5 = 45 ohms

Regards
Vikas TU
4 years ago
Dear student 
R1 = 5 ohm 
l1 = l 
l2 = 2l
When wire is stretched V1 = V2 
A1l1 = A2l2 
A2 = A1/2
R1/R2 = 1/4 
R2 = 4R1